OpenStudy (anonymous):

Write an equation of the line that passes through (8,3) and (4,3) y=3 y=4x+3 y=4 y=3x

OpenStudy (anonymous):

First, find the slope. Do you know how to find the slope of two points?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

To find the slope of two points, you do \[\frac{ y _{2}-y _{1} }{x _{2}-x _{1} }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y2 is 3, y1 is 3, and x2 is 4, and x1 is 8

OpenStudy (anonymous):

3-3 4-8?

OpenStudy (anonymous):

yes, so that means the slope is 0/-4 which is 0.

OpenStudy (anonymous):

now plug that into the point slope form equation which is y-y1=m(x-x1)

OpenStudy (anonymous):

where does division come from

OpenStudy (anonymous):

very confusing

OpenStudy (anonymous):

the reason you divide is because to find the slope, that is how you get it. The explanation is long, but if you want it I can tell you

OpenStudy (anonymous):

so its y=3?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

ok ty so much

OpenStudy (anonymous):

Another way to solve this is looking at the two points, and seeing they both have the same y coordinate of 3. That means that y=3

OpenStudy (anonymous):

can you help me with 1 more

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

what does the slope for this relation represent x(years since 1991) 2,5,7,7 y(book sales billions$) 18,21,23,25

OpenStudy (anonymous):

its kind of like a little chart

OpenStudy (anonymous):

Well, if the chart lines up so that 2 and 18 go together and 5 and 21 go together and so on, then it is just the points (2, 18) (5, 21) etc

OpenStudy (anonymous):

ect?

OpenStudy (anonymous):

it means and so forth

OpenStudy (anonymous):

just wait, ill be back in like a minute

OpenStudy (anonymous):

so thats the answer

OpenStudy (anonymous):

2,18 5,21

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Im back

OpenStudy (anonymous):

im guessing can my question a couple of seconds ago

OpenStudy (anonymous):

read my*

OpenStudy (anonymous):

still here?

OpenStudy (anonymous):

did you get the last x coordinate wrong? should it be 8 or is it 7

OpenStudy (anonymous):

its 8 oh

OpenStudy (anonymous):

Is it possible to attach a picture of the question?

OpenStudy (anonymous):

no but let me write it the best i can humanly do it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

X(Years Since 1991) 2,5,7,8 Y(Book Sales Billions$)18,21,23,24 What does the slope for this relation REPRESENT

OpenStudy (anonymous):

Oh, you don't have to find the slope, you just need to know why it is important

OpenStudy (anonymous):

if it says that 2 years after 1991 you sell 18 billion books and 5 years after 1991 you sell 21 billion books, what does that tell you

OpenStudy (anonymous):

1991 to 1996 21 billoin books sold

OpenStudy (anonymous):

Sort of,

OpenStudy (anonymous):

i mean 3 billion from 1991 to 1996

OpenStudy (anonymous):

it says that after 2 years you sell 18 billion and after 5 years you sell 21 billion, that means that the slope shows the increase or decrease in sales for those years

OpenStudy (anonymous):

increase!!!!!!!

OpenStudy (anonymous):

oh wait its a decrease

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

No, actually it is a decrease because in two years you sell 18 billion, but then in the next 3 years, you only sell 3 billion. Then in the next 2 years you sell 2 billion and the next year you sell 1 billion

OpenStudy (anonymous):

That means for the first you years, you sell 9 billion a year, but then after 8 years, you only sell 1 billion a year

OpenStudy (anonymous):

*the first two years

OpenStudy (anonymous):

so the relation represents a decrease or somthing

OpenStudy (anonymous):

Yes, the slope shows the decrease in sales.

OpenStudy (anonymous):

awsome stuff ty alot

OpenStudy (anonymous):

yw